x1234y56912 Find the gradient of the line of best fit of the data.


1
Expert's answer
2021-09-01T11:58:26-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 x & y \\ \hline
 1 & 5 \\ \hline
 2 & 6 \\ \hline
 3 & 9 \\
 \hdashline
 4 & 12 
\end{array}"

"\bar{x}=\dfrac{\sum_ix_i}{n}=\dfrac{1+2+3+4}{4}=2.5"

"\bar{y}=\dfrac{\sum_iy_i}{n}=\dfrac{5+6+9+12}{4}=8"

"S_{xx}=\sum_i(x_i-\bar{x})^2=\sum_ix_i^2-n\cdot\bar{x}^2=5"

"S_{xy}=\sum_i(x_i-\bar{x})(y_i-\bar{y})=\sum_ix_iy_i-n\cdot\bar{x}\bar{y}=12"

"S_{yy}=\sum_i(y_i-\bar{y})^2=\sum_iy_i^2-n\cdot\bar{y}^2=30"

"gradient=\dfrac{S_{xy}}{S_{xx}}=\dfrac{12}{5}=2.4"